How would you graph
y = ln x + 4
I know the asypmtote would be @ vertical -4
but in not sure how to make the T chart, what to plug in for x, etc.

Answer by jsmallt9(3758)

First, please put parentheses around the arguments of logarithms. What you posted looks like
y = ln(x) + 4
It is only because you mentioned a vertical asymptote of -4 that I know that the equation is really:
y = ln(x + 4)
These two equations are different and have different graphs. Tutors are more likely to help if the problem is clear.

Since the base of ln is "e", it is difficult to find most ln's without a calculator. So all you have to do is:
  1. Pick a value for x, (Note: Since arguments to logarithms must be positive, you must pick a number greater thatn -4!)
  2. Add 4 to the value of x
  3. Use your calculator to find the ln of the result of step #2. This will be the value for y.
  4. Graph this point.
  5. Repeat all of the above until you can see how the graph goes.
  6. Draw a smooth curve through the points you've graphes.

If you know what the graph of
y = ln(x)
looks like and if you understand about transformations of graphs, then you could graph
y = ln(x + 4)
very quickly because its graph is exactly the same as the graph of
y = ln(x)
except that it is moved to the left 4 units.